OpenStudy (anonymous):

Give me the definitions of these Sidereal period Sidereal day Synodic Period and the equations for finding those of the solar system planets

OpenStudy (anonymous):

look online that might help u

OpenStudy (anonymous):

I looked it up but not sufficient for me

OpenStudy (anonymous):

sidereal period, the time required for a celestial body within the solar system to complete one revolution with respect to the fixed stars—i.e., as observed from some fixed point outside the system. The sidereal period of a planet can be calculated if its synodic period (the time for it to return to the same position relative to the Sun and Earth) is known; the sidereal period of the Moon or an artificial satellite of Earth is the time needed for it to return to the same position against the background of stars. See also synodic period.

OpenStudy (anonymous):

Sidereal Day The length of time which passes between a given "fixed" star in the sky crossing a given projected meridian (line of longitude). The sidereal day is 23 h 56 m 4.1 s, slightly shorter than the solar day because the Earth Eric Weisstein's World of Astronomy's orbital motion about the Sun Eric Weisstein's World of Astronomy means the Earth Eric Weisstein's World of Astronomy has to rotate slightly more than one turn with respect to the "fixed" stars in order to reach the same Earth-Sun orientation. Another way of thinking about the difference is that it amounts to 1/365.2425th of a day per day, since even if the Earth Eric Weisstein's World of Astronomy did not spin on its axis at all, the Sun Eric Weisstein's World of Astronomy would appear to make one rotation around the Earth Eric Weisstein's World of Astronomy as the Earth Eric Weisstein's World of Astronomy completed a single orbit (which takes one year).

OpenStudy (anonymous):

synodic period, the time required for a body within the solar system, such as a planet, the Moon, or an artificial Earth satellite, to return to the same or approximately the same position relative to the Sun as seen by an observer on the Earth. The Moon’s synodic period is the time between successive recurrences of the same phase; e.g., between full moon and full moon. The synodic period of a planet is the time required for the Earth to overtake it as both go around the Sun—or, in the case of fast-moving Mercury or Venus, for the planet in question to overtake the Earth. The synodic period of an artificial satellite of the Earth is measured between its conjunctions (closest apparent approaches) with the Sun. See also sidereal period. LINKS

OpenStudy (anonymous):

I really don't know - but this page has a bunch of formulas on it: http://en.wikipedia.org/wiki/Orbital_period

OpenStudy (anonymous):

I want equations

OpenStudy (anonymous):

N = R^{\ast} \cdot f_p \cdot n_e \cdot f_{\ell} \cdot f_i \cdot f_c \cdot L \! idk if this is right tho

OpenStudy (anonymous):

What is this

OpenStudy (anonymous):

I'm happy if I helped you and if I didn't im srry i gotta go